y"+3y=0 , y(0)=0 and y'(3)=0


1
Expert's answer
2021-05-12T01:58:00-0400

Given, y''+3y=0,y(0)=0 and y'(3)=0


"\Rightarrow y''+3y=0"


"(D^2+3)y=0"


So Auxiliary equation-

"m^2+3=0\Rightarrow m=\pm \sqrt{3}i"


Solution of the equation is-


"y=c_1cos\sqrt{3}x+c_2sin\sqrt{3}x~~~~~-(1)"


At x=0,y=0


"0=c_1cos0+c_2sin 0\Rightarrow c_1=0"


Differentiating equation (1) w.r.t. x-


"y'=-\sqrt{3}c_1sin\sqrt{3}x+\sqrt{3}c_2cos\sqrt{3}x"


Also, "y'(3)=0"


"\Rightarrow 0=-\sqrt{3}c_1sin3\sqrt{3}+\sqrt{3}c_2 cos3\sqrt{3}"


As "c_1=0\Rightarrow c_2=0"


Solution of equation is

y=0